OpenStudy (anonymous):

*Correct Me :D !! Titus deposited $1,508.49 in a savings account that earns 4.2% simple interest. What will Titus’s account balance be in 8 months? $2,015.34<----- $1,571.85 $1,508.49 $1,550.73 --------- MY WORK **** 1508.49*4.2%*8+1508.49=2015.34264 ?

OpenStudy (anonymous):

You're decimal is off for the interest rate. Should be multiplying by .042

OpenStudy (anonymous):

1508.49*.042+1508.49 = ?

jimthompson5910 (jim_thompson5910):

also keep in mind that it's in the account for 8 months, not a full year

jimthompson5910 (jim_thompson5910):

so t = 8/12 and not t = 1

jimthompson5910 (jim_thompson5910):

oh and it's definitely not t = 8 either, unless you wanted the money to stay in for 8 years